OpenStudy (anonymous):

Given that sin(a)=k, and that 90

OpenStudy (freckles):



OpenStudy (freckles):

you can find the adjacent side by using Pythagorean theorem

OpenStudy (anonymous):

omg thanks, so cos(a)=sqrt(1-k^2) and tan(a)=(k(sqrt(1-k^2)))/(1-k^2)?

OpenStudy (freckles):

well I would be careful with signs k^2+adj^2=1^2 adj=plus or minus sqrt(1-k^2) but since the adj is an x value and it is to the left of 0 then we choose adj=-sqrt(1-k^2)

OpenStudy (anonymous):

Oh yeah oops, forgot about that. Thanks so much though!!

OpenStudy (freckles):

np